# Dynamic-SUPERB Phase-2: A Collaboratively Expanding Benchmark for Measuring the Capabilities of Spoken Language Models with 180 Tasks

### v1

@@ Appendix C Task-level evaluation results
| Domestic Environment Sound Event Detection | NAR | 91.39% | 99.44% | 99.72% | 99.72% | 99.17% | 99.72% | 99.72% | 100.00% | 99.72% | |
| Event Based F1 | 0.0000 | 0.6000 | 0.0000 | 0.0000 | 0.0000 | 0.0100 | 0.0000 | 0.0000 | 0.0000 | ||
| Environmental Sound Classification (UrbanSound8K UrbanNoises) | LLM-C | 10.00% | 3.26% | 1.40% | 0.00% | 39.53% | 4.88% | 10.47% | 0.00% | 0.70% | |
| HEAR Environmental Sound Classification (ESC-50) | LLM-C | 4.80% | 25.30% | 61.00% | 73.40% | 68.80% | 43.90% | 9.80% | 0.40% | 35.30% | |
| Environmental Sound Classification (ESC50 Natural Soundscapes And Water Sound) | LLM-C | 3.00% | 7.50% | 6.50% | 1.00% | 78.00% | 5.00% | 7.00% | 0.00% | 7.00% | |
| Environmental Sound Classification (ESC50 Exterior And UrbanNoises) | LLM-C | 4.50% | 14.50% | 0.50% | 0.50% | 77.00% | 25.00% | 6.50% | 0.00% | 6.50% | |
| Environmental Sound Classification (ESC50 Interior And Domestic Sounds) | LLM-C | 4.50% | 3.50% | 0.50% | 0.00% | 59.00% | 2.50% | 4.00% | 0.00% | 0.00% | |
| Instrument Source Classification | LLM-C | 7.22% | 28.33% | 37.33% | 43.56% | 32.67% | 35.22% | 30.89% | 38.00% | 35.89% | |
| Instrument Classification (Beijing Opera Instrument) | LLM-C | 3.39% | 10.59% | 6.36% | 23.73% | 30.93% | 28.39% | 22.03% | 25.85% | 10.17% | |
| Instrument Combination Recognition | LLM-C | 4.20% | 13.51% | 17.72% | 24.02% | 25.83% | 24.92% | 20.42% | 19.52% | 8.11% | |
| Instrument Classification (Nsynth) | LLM-C | 1.23% | 9.56% | 13.97% | 14.58% | 17.89% | 54.41% | 11.89% | 6.99% | 10.17% | |
| HEAR Percussion Instruments Classification | LLM-C | 3.39% | 4.24% | 10.17% | 20.76% | 16.10% | 27.12% | 24.15% | 22.03% | 11.44% | |
| HEAR Percussion Instruments Stroke Classification | LLM-C | 0.10% | 3.80% | 7.60% | 4.90% | 11.50% | 10.80% | 13.70% | 10.30% | 0.10% | |
| MARBLE Instrument Classification | LLM-C | 1.00% | 1.00% | 0.90% | 0.00% | 9.70% | 3.30% | 1.10% | 1.00% | 0.60% | |
| Animal Classification | LLM-C | 9.25% | 40.50% | 58.25% | 67.00% | 75.75% | 34.00% | 12.00% | 4.25% | 15.25% | |
| Cornell Birdcall Identification | LLM-C | 0.00% | 0.00% | 0.00% | 3.33% | 10.00% | 10.00% | 13.33% | 3.33% | 0.00% | |
| Cat Emotion Classification | LLM-C | 6.00% | 0.00% | 2.00% | 4.00% | 6.00% | 14.00% | 8.00% | 2.00% | 14.00% | |
| HEAR Beehive States Classification | LLM-C | 43.00% | 35.00% | 54.00% | 42.50% | 19.00% | 42.50% | 37.00% | 37.50% | 19.00% | |
| Environmental Sound Classification | LLM-C | 14.50% | 8.00% | 14.00% | 3.00% | 82.00% | 36.00% | 5.00% | 0.00% | 2.00% | |
| Bird Sound Detection | LLM-C | 28.50% | 43.00% | 74.50% | 75.00% | 78.50% | 79.00% | 75.00% | 40.50% | 33.00% | |
| Music Genre Classification | LLM-C | 10.71% | 1.79% | 9.82% | 16.07% | 8.04% | 15.18% | 9.82% | 9.82% | 4.46% | |
| HEAR Music Genre Classification | LLM-C | 7.30% | 9.70% | 37.50% | 42.50% | 29.00 % | 29.40 % | 9.50% | 26.90% | 7.00% | |
| MARBLE Music Tagging (MTGTop50) | LLM-C | 0.00% | 0.10% | 0.00% | 0.00% | 1.40% | 0.20% | 0.00% | 0.60% | 0.10% | |
| MARBLE Music Tagging (MagnaTagATune) | LLM-C | 5.50% | 1.00% | 3.50% | 9.00% | 4.50% | 1.50% | 1.00% | 3.50% | 7.50% | |
| MARBLE Genre Classification (MTG-Genre) | LLM-C | 0.60% | 2.10% | 0.00% | 0.00% | 3.70% | 0.80% | 0.80% | 0.80% | 0.30% | |
| MARBLE Genre Classification (GTZAN) | LLM-C | 28.50% | 8.00% | 32.00% | 61.50% | 63.50% | 68.50% | 22.50% | 17.00% | 5.00% | |
| Sound Effect Detection | LLM-C | 0.81% | 3.00% | 4.56% | 10.06% | 7.81% | 13.69% | 7.38% | 7.31% | 3.31% | |
| Audio Duration Prediction | NAR | 33.00% | 37.00% | 10.50% | 55.00% | 0.00% | 0.00% | 0.50% | 46.50% | 14.00% | |
| MSE | 13.3657 | 39.1111 | 31652.7877 | 2216.7111 | 28.985 | 62.515 | 3581.0352 | 1527.785 | 62.7674 | ||
| HEAR Music Speech Classification | LLM-C | 63.28% | 34.38% | 74.22% | 75.78% | 96.88% | 75.00% | 60.94% | 53.13% | 82.81% | |
| Speech Detection (LJSpeech) | LLM-C | 54.00% | 44.50% | 60.50% | 74.00% | 57.00% | 45.00% | 53.00% | 49.00% | 55.50% | |
| Speech Detection (LibriSpeech-TestOther) | LLM-C | 52.00% | 31.50% | 46.00% | 66.50% | 44.50% | 55.50% | 37.00% | 45.50% | 36.00% | |
| Speech Detection (LibriSpeech-TestClean) | LLM-C | 48.00% | 29.00% | 46.00% | 66.00% | 45.50% | 56.50% | 35.00% | 46.50% | 38.50% | |
| Phonological Feature Classification (VoxAngeles-ConsonantPlaceOfArticulation) | LLM-C | 25.77% | 1.37% | 1.88% | 1.37% | 1.54% | 3.24% | 1.88% | 1.54% | 0.51% | |
| Phonological Feature Classification (VoxAngeles-MannerOfArticulation) | LLM-C | 17.42% | 9.36% | 2.06% | 1.50% | 5.62% | 6.74% | 6.84% | 7.96% | 8.99% | |
| Phonological Feature Classification (VoxAngeles-Phone) | LLM-C | 5.73% | 0.09% | 1.76% | - | 3.60% | 3.97% | 0.00% | 0.00% | 0.09% | |
| Phonological Feature Classification (VoxAngeles-VowelFrontnes) | LLM-C | 50.71% | 17.52% | 41.14% | 38.09% | 48.47% | 57.23% | 36.46% | 42.97% | 10.39% | |
| Phonological Feature Classification (VoxAngeles-VowelHeight) | LLM-C | 24.44% | 17.11% | 31.77% | 35.23% | 29.53% | 36.86% | 37.88% | 39.10% | 24.44% | |
| Phonological Feature Classification (VoxAngeles-VowelRoundedness) | LLM-C | 43.38% | 21.18% | 27.29% | 61.30% | 46.44% | 69.65% | 38.90% | 21.79% | 18.94% | |
| L2 English Prosodic (speechocean762-Ranking) | LLM-C | 26.39% | 35.00% | 46.67% | 52.50% | 34.44% | 51.11% | 34.44% | 49.72% | 42.22% | |
| NAR | 39.43% | 94.75% | 0.27% | 1.21% | 71.47% | 3.10% | 11.98% | 94.62% | 82.91% | ||
| PCC | 0.0435 | -0.1747 | 0.0427 | 0.0775 | 0.1446 | 0.0201 | 0.0461 | -0.1535 | 0.0973 | ||
| L2 English Fluency (speechocean762-Ranking) | LLM-C | 31.39% | 21.94% | 50.00% | 50.00% | 40.28% | 50.56% | 27.78% | 49.44% | 45.00% | |
| NAR | 14.67% | 88.16% | 0.00% | 0.67% | 8.75% | 0.81% | 41.86% | 63.53% | 55.72% | ||
| PCC | 0.0055 | 0.0332 | 0.0292 | 0.0183 | -0.0422 | -0.0858 | 0.0222 | 0.0505 | 0.0532 | ||
| Prosody Naturalness (ProsAudit-Lexical) | LLM-C | 49.81% | 32.82% | 48.26% | 47.10% | 21.62% | 51.74% | 54.83% | 47.10% | 5.41% | |
| Prosody Naturalness (ProsAudit-Protosyntax) | LLM-C | 53.44% | 31.68% | 46.56% | 46.18% | 25.95% | 50.76% | 43.51% | 49.24% | 4.20% | |
| SuperbSD | NAR | 93.00% | 99.50% | 100.00% | 100.00% | 99.50% | 98.00% | 97.50% | 100.00% | 60.50% | |
| diarization error rate | 74.81% | 74.62% | - | - | 44.26% | 93.29% | 90.53% | - | 74.12% | ||
| SuperbSV | LLM-C | 53.00% | 40.00% | 65.50% | 93.50% | 47.50% | 55.50% | 66.50% | 49.00% | 33.50% | |
| HEAR Speaker Count Identification | NAR | 30.40% | 42.10% | 35.90% | 42.50% | 3.50% | 0.40% | 22.00% | 23.20% | 26.20% | |
| Acc | 17.53% | 16.93% | 13.26% | 17.22% | 10.26% | 13.76% | 5.51% | 13.80% | 17.48% | ||
| Speaker Counting | LLM-C | 18.50% | 6.00% | 14.50% | 13.50% | 17.50% | 26.50% | 14.00% | 4.00% | 12.00% | |
| Multi Speaker Detection | LLM-C | 46.50% | 23.50% | 36.00% | 30.50% | 57.00% | 51.00% | 32.50% | 39.00% | 48.50% | |
| Speaker Verification (LibriSpeech-TestClean) | LLM-C | 48.00% | 37.50% | 45.00% | 54.50% | 43.50% | 52.00% | 19.00% | 24.00% | 13.00% | |
| Task | Metric | Whisper- | LTU-AS | SALMONN | SALMONN | Qwen- | Qwen2- | WavLLM | MU- | GAMA |
| LLaMA | 7B | 13B | Audio | Audio | LLaMA | |||||
| SuperbST | sacreBLEU | 17.3720 | 0.1070 | 18.7884 | 16.7835 | 6.9846 | 23.3458 | 21.7423 | 0.0566 | 0.0219 |
@@
| Third Tone Sandhi Recognition | NAR | 59.38% | 93.75% | 100.00% | 87.50% | 84.38% | 84.38% | 28.12% | 96.88% | 100.00% |
| IoU | 0.2179 | 0.5 | 0 | 0.75 | 0 | 0 | 0.1304 | 0 | 0 | |
| Accent Classification | LLM-C | 17.50% | 5.50% | 3.00% | 4.50% | 26.50% | 14.00% | 7.00% | 27.00% | 4.00% |
| Stress Detection | LLM-C | 15.50% | 3.00% | 2.00% | 13.00% | 16.50% | 23.50% | 25.50% | 0.00% | 1.00% |
| Phoneme Segmentation Counting | NAR | 19.19% | 57.24% | 41.88% | 12.56% | 0.41% | 16.11% | 15.04% | 41.47% | 18.79% |
| Acc | 13.20% | 1.69% | 15.81% | 20.04% | 1.77% | 6.60% | 10.11% | 9.14% | 3.80% | |
| Abs Diff | 41.236 | 3759.3878 | 5.4857 | 4.9331 | 3.1286 | 2.8706 | 45754.9952 | 5.4242 | 2.8694 | |
| Phoneme Segmentation Counting | NAR | 16.29% | 48.64% | 39.09% | 11.62% | 5.99% | 1.73% | 0.50% | 2.67% | 72.22% |
| Acc | 13.56% | 11.77% | 11.84% | 8.25% | 5.42% | 13.46% | 9.55% | 0.47% | 2.29% | |
| Abs Diff | 2.5378 | 1012.8103 | 6.5285 | 50.7724 | 7.1857 | 3.0622 | 7.9109 | 15.6777 | 5.0491 | |
| SUPERB PR | Phoneme Error Rate | 100.12% | 102.75% | 25.36% | 24.60% | 100.58% | 100.96% | 99.99% | 110.27% | 100.37% |

### v2

@@ Appendix C Task-level evaluation results
| Domestic Environment Sound EventDetection (DESED-PublicEval) | Event-based F1 | 0 | 0.6 | 0 | 0 | 0 | 0.01 | 0 | 0 | 0 |
| Emergency Traffic Detection (Large-Scale-Audio-dataset) | LLM-C | 39.00% | 34.00% | 55.50% | 7.25% | 65.50% | 74.50% | 17.50% | 43.50% | 15.83% |
| Environmental Sound Classification (ESC50-Exterior And Urban Noises) | LLM-C | 4.50% | 14.50% | 0.50% | 0.50% | 77.00% | 25.00% | 6.50% | 0.00% | 6.50% |
| Environmental Sound Classification (ESC50-Interior And Domestic Sounds) | LLM-C | 4.50% | 3.50% | 0.50% | 0.00% | 59.00% | 2.50% | 4.00% | 0.00% | 0.00% |
| Environmental Sound Classification (ESC50-Natural Soundscapes And Water Sounds) | LLM-C | 3.00% | 7.50% | 6.50% | 1.00% | 78.00% | 5.00% | 7.00% | 0.00% | 7.00% |
| Environmental Sound Classification (UrbanSound8K-Urban Noises) | LLM-C | 10.00% | 3.26% | 1.40% | 0.00% | 39.53% | 4.88% | 10.47% | 0.00% | 0.70% |
| Environment Recognition (ESC50) | LLM-C | 7.02% | 38.60% | 42.11% | 19.30% | 56.14% | 63.16% | 40.35% | 10.53% | 31.58% |
| HEAR Environmental Sound Classification (ESC50) | LLM-C | 4.80% | 25.30% | 61.00% | 73.40% | 68.80% | 43.90% | 9.80% | 0.40% | 35.30% |
| HEAR Vocal Imitation Classification (Vocal Imitations) | LLM-C | 15.17% | 5.50% | 1.17% | 1.83% | 18.67% | 11.17% | 17.50% | 3.83% | 5.83% |
| Audio Spatial Distance Prediction (Spatial LibriSpeech) | NAR | 50.50% | 59.00% | 62.50% | 100.00% | 85.50% | 58.00% | 44.00% | 97.00% | 71.00% |
| Audio Spatial Distance Prediction (Spatial LibriSpeech) | MEDAE | 0.6651 | 0.933 | 0.449 | - | 1.5011 | 0.6915 | 0.9011 | 0.9212 | 0.5 |
| How Far Are You (3DSpeaker) | LLM-C | 26.00% | 13.00% | 4.50% | 8.00% | 30.00% | 26.00% | 32.00% | 18.00% | 2.50% |
| Sound Position Prediction | NAR | 81.25% | 100.00% | 100.00% | 100.00% | 100.00% | 25.00% | 93.75% | 100.00% | 100.00% |
| Sound Position Prediction | Angle Diff | 1.4006 | - | - | - | - | 1.1063 | 0.8681 | - | - |
| Dialogue Emotion Classification (Daily Talk) | LLM-C | 33.50% | 16.50% | 27.00% | 11.50% | 33.00% | 54.00% | 18.50% | 19.00% | 33.00% |
| Emoji Grounded Speech Emotion Recognition (RAVDESS) | LLM-C | 1.10% | 0.00% | 0.00% | 0.10% | 1.30% | 1.20% | 0.10% | 1.90% | 48.50% |
| Emotion Change Detection (Ravdess) | LLM-C | 0.08% | 0.42% | 0.58% | 0.96% | 0.04% | 2.96% | 0.08% | 0.58% | 0.04% |
| Emotion Recognition (MELD) | LLM-C | 22.50% | 14.50% | 19.00% | 9.50% | 43.50% | 36.50% | 40.50% | 0.50% | 2.50% |
| HEAR Emotion Recognition (CREMAD) | LLM-C | 6.70% | 9.30% | 18.80% | 18.10% | 62.50% | 61.10% | 25.60% | 12.40% | 13.20% |
| SUPERB Emotion Recognition (RAVDESS) | LLM-C | 12.50% | 25.83% | 12.08% | 12.50% | 70.42% | 75.83% | 12.08% | 12.92% | 0.00% |
| Covid19 Cough Audio Classification (CoughVid) | LLM-C | 0.93% | 1.03% | 0.00% | 0.00% | 4.12% | 0.72% | 4.22% | 21.91% | 0.21% |
| Environmental Sound Classification (ESC50-Human And Non Speech Sounds) | LLM-C | 17.50% | 11.00% | 8.50% | 3.00% | 82.00% | 29.00% | 16.50% | 0.00% | 2.00% |
| Human Non Speech Sound Recognition (Nonspeech7k, CommonVoice) | LLM-C | 2.86% | 5.00% | 44.29% | 30.71% | 27.14% | 23.57% | 14.29% | 17.14% | 32.86% |
| Human Screaming Detection (Environmentdb) | LLM-C | 52.50% | 25.00% | 52.50% | 65.00% | 62.50% | 87.50% | 47.50% | 42.50% | 60.00% |
| Vocal Sound Recognition (VocalSound) | LLM-C | 29.86% | 1.53% | 11.53% | 4.31% | 75.56% | 30.97% | 14.44% | 0.28% | 2.78% |
| Phoneme Segment Counting (Librispeech-words) | NAR | 16.29% | 48.64% | 39.09% | 11.62% | 5.99% | 1.73% | 0.50% | 2.67% | 72.22% |
| Phoneme Segment Counting (Librispeech-words) | ACC | 13.56% | 11.77% | 11.84% | 8.25% | 5.42% | 13.46% | 9.55% | 0.47% | 2.29% |
| Phoneme Segment Counting (Librispeech-words) | Abs Diff | 2.5378 | 1012.8103 | 6.5285 | 50.7724 | 7.1857 | 3.0622 | 7.9109 | 15.6777 | 5.0491 |
| Phone Segment Counting (VoxAngeles) | NAR | 19.19% | 57.24% | 41.88% | 12.56% | 0.41% | 16.11% | 15.04% | 41.47% | 18.79% |
| Phone Segment Counting (VoxAngeles) | ACC | 13.20% | 1.69% | 15.81% | 20.04% | 1.77% | 6.60% | 10.11% | 9.14% | 3.80% |
| Phone Segment Counting (VoxAngeles) | Abs Diff | 41.236 | 3759.3878 | 5.4857 | 4.9331 | 3.1286 | 2.8706 | 45754.9952 | 5.4242 | 2.8694 |
| SUPERB Phoneme Recognition (LibriSpeech-TestClean) | PER | 100.12% | 102.75% | 25.36% | 24.60% | 100.58% | 100.96% | 99.99% | 110.27% | 100.37% |
| Phoneme Recognition (LibriSpeech-TestOther) | PER | 100.17% | 101.08% | 22.79% | 22.91% | 100.62% | 100.88% | 99.99% | 111.18% | 100.61% |
| Phonological Feature Classification (VoxAngeles-Consonant Place Of Articulation) | LLM-C | 25.77% | 1.37% | 1.88% | 1.37% | 1.54% | 3.24% | 1.88% | 1.54% | 0.51% |
| Phonological Feature Classification (VoxAngeles-Manner Of Articulation) | LLM-C | 17.42% | 9.36% | 2.06% | 1.50% | 5.62% | 6.74% | 6.84% | 7.96% | 8.99% |
| Phonological Feature Classification (VoxAngeles-Phone) | LLM-C | 5.73% | 0.09% | 1.76% | - | 3.60% | 3.97% | 0.00% | 0.00% | 0.09% |
| Phonological Feature Classification (VoxAngeles-Vowel Frontness) | LLM-C | 50.71% | 17.52% | 41.14% | 38.09% | 48.47% | 57.23% | 36.46% | 42.97% | 10.39% |
| Phonological Feature Classification (VoxAngeles-Vowel Height) | LLM-C | 24.44% | 17.11% | 31.77% | 35.23% | 29.53% | 36.86% | 37.88% | 39.10% | 24.44% |
| Phonological Feature Classification (VoxAngeles-Vowel Roundedness) | LLM-C | 43.38% | 21.18% | 27.29% | 61.30% | 46.44% | 69.65% | 38.90% | 21.79% | 18.94% |
| Heteronym Differentiation (HeteronymEn) | LLM-C | 55.00% | 26.00% | 37.00% | 45.00% | 52.00% | 44.00% | 51.00% | 30.00% | 20.00% |
| L2 English Accuracy (speechocean762-Ranking) | LLM-C | 24.72% | 33.89% | 50.00% | 49.72% | 34.44% | 50.00% | 41.67% | 48.06% | 45.56% |
| L2 English Accuracy (speechocean762-scoring) | NAR | 15.44% | 95.03% | 0.40% | 0.27% | 37.72% | 0.94% | 9.53% | 90.47% | 50.07% |
| L2 English Accuracy (speechocean762-scoring) | PCC | 0.0185 | -0.183 | 0.0633 | 0.0438 | 0.0293 | -0.0159 | 0.0439 | 0.0727 | 0.0151 |
| Multi-lingual Pronunciation Similarity (VoxAngeles) | LLM-C | 38.40% | 13.50% | 47.20% | 48.50% | 23.70% | 44.70% | 48.00% | 25.10% | 40.60% |
| Accent Classification (Accentdb Extended) | LLM-C | 17.50% | 5.50% | 3.00% | 4.50% | 26.50% | 14.00% | 7.00% | 27.00% | 4.00% |
| Stress Detection (MIRSD) | LLM-C | 15.50% | 3.00% | 2.00% | 13.00% | 16.50% | 23.50% | 25.50% | 0.00% | 1.00% |
@@
| Third Tone Sandhi Recognition (NCCU Corpus of Spoken Taiwan Mandarin) | NAR | 59.38% | 93.75% | 100.00% | 87.50% | 84.38% | 84.38% | 28.12% | 96.88% | 100.00% |
| Third Tone Sandhi Recognition (NCCU Corpus of Spoken Taiwan Mandarin) | IoU | 0.2179 | 0.5 | 0 | 0.75 | 0 | 0 | 0.1304 | 0 | 0 |
| L2 English Fluency (speechocean762-Ranking) | LLM-C | 31.39% | 21.94% | 50.00% | 50.00% | 40.28% | 50.56% | 27.78% | 49.44% | 45.00% |
| L2 English Fluency (speechocean762-Scoring) | NAR | 14.67% | 88.16% | 0.00% | 0.67% | 8.75% | 0.81% | 41.86% | 63.53% | 55.72% |
| L2 English Fluency (speechocean762-Scoring) | PCC | 0.0055 | 0.0332 | 0.0292 | 0.0183 | -0.0422 | -0.0858 | 0.0222 | 0.0505 | 0.0532 |
| L2 English Prosodic (speechocean762-Ranking) | LLM-C | 26.39% | 35.00% | 46.67% | 52.50% | 34.44% | 51.11% | 34.44% | 49.72% | 42.22% |
| L2 English Prosodic (speechocean762-Scoring) | NAR | 39.43% | 94.75% | 0.27% | 1.21% | 71.47% | 3.10% | 11.98% | 94.62% | 82.91% |
| L2 English Prosodic (speechocean762-Scoring) | PCC | 0.0435 | -0.1747 | 0.0427 | 0.0775 | 0.1446 | 0.0201 | 0.0461 | -0.1535 | 0.0973 |
| Prosody Naturalness (ProsAudit-Lexical) | LLM-C | 49.81% | 32.82% | 48.26% | 47.10% | 21.62% | 51.74% | 54.83% | 47.10% | 5.41% |
| Prosody Naturalness (ProsAudit-Protosyntax) | LLM-C | 53.44% | 31.68% | 46.56% | 46.18% | 25.95% | 50.76% | 43.51% | 49.24% | 4.20% |